OpenStudy (anonymous):

Find the equation of the line (1, -2) and (5, -4) in the form y=mx +b

OpenStudy (anonymous):

\[y=-\frac{ 1 }{ 2 }x -\frac{ 3 }{ 2 }\]

OpenStudy (johnweldon1993):

you use this formula (y2-y1)/(x2-x1) and you have 2 points...which is all you need (-4 + 2)/(5-1) -2/4 simplified is -1/2 this is your slope now apply this to your point slope formula y - y1 = m(x - x1) m is your slope pick one of the 2 points you have...and put the y where you see y1...and put the x where you see x1 so y - (-2) = -1/2(x - 1) y + 2 = -1/2x + 1/2 subtract 2 from both sides y = -1/2x + (1/2 - 2) find lcd of 1/2 and 2 which means 1/2 - 4/2 which becomes -3/2 so now you have y = -1/2x - 3/2